OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative. I don't get this in word problem form

OpenStudy (anonymous):

what's the derivative of the given s(t) with respect to t?

OpenStudy (anonymous):

s(t) represents the distance in terms of variable t. instantaneous velocity ( velocity ) is - as we know - is the rate of change of distance . In math terms - rate of change is given by derivative. so we need to have derivative of s(t) to have velocity. instantaneous velocity is the velocity at a particular instant , here he's asking for velocity at t = 1 :)

OpenStudy (anonymous):

find the derivative of s(t). substitute t = 1 in the result and there you will have answer to the question.

OpenStudy (anonymous):

so do you just plug in 1 for x? I'm confused since there's no exponent

OpenStudy (anonymous):

I just dont get how to find the derivative since theres no exponent because I only knew how to do the power rule

OpenStudy (anonymous):

when its given \[x^2\] you know its exponent is 2 right? that's how when it's just x. you consider it \[x^1\] -> exponent is 1. :)

OpenStudy (anonymous):

OH IM DUMB, thank you! Let me do this real quick

OpenStudy (anonymous):

no problem, it happens at times :)

OpenStudy (anonymous):

Okay, so since it's 1, nothing really changes, right? I got -17 as the answer but I know thats wrong.

OpenStudy (anonymous):

derivative of constant is 0. FYI

OpenStudy (anonymous):

Ugh, I'm really sorry but I don't understand that.. Could you explain?

OpenStudy (freckles):

do you know this definition: \[\lim_{h \rightarrow 0}\frac{s(t+h)-s(t)}{h}=s'(t)\]

OpenStudy (anonymous):

I don't know how to use it, no

OpenStudy (freckles):

actually you have a line just fine the slope of your line

OpenStudy (freckles):

what is the slope of s(t)=-9t-8?

OpenStudy (freckles):

for example the slope of y=mx+b is m so the slope of y=-9x-8 is ?

OpenStudy (anonymous):

hmm, derivative with respect to t means how equation changes with respect to variable t. so when we say \[t^1\] now for t= 1 -> we have t^1 values = 1 for t =2 -> we havr t^1 values = 2 for t =3 -> we have t^1 values = 3 so change is 2-1 = 3-2 = 1{ constant value }

OpenStudy (anonymous):

-9

OpenStudy (freckles):

that is your answer

OpenStudy (freckles):

derivative=slope

OpenStudy (anonymous):

really?? That's all it is?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

@SandeepReddy , oh I understand now!

OpenStudy (anonymous):

Thank you both so much!!!(:

OpenStudy (anonymous):

pleasure is ours :)